OpenStudy (anonymous):

How can you solve the equation y- 6.5 = 7.1

OpenStudy (anonymous):

Add 6.5 both the sides to find y..

OpenStudy (anonymous):

your job is to isolate y, so its all by itself on one side you can go about doing so if you added 6.5 to both sides, which would remove the 6.5 on the left side \[y-6.5=7.1\] \[y-6.5+6.5=7.1+6.5\] \[y=7.1+6.5\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

just remember for these types of problems, if you ever multiply, divide, add, or subtract to once side. you have to do it for the other side